# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ A.3 Concentration of Empirical FID (4)
Recall that for any generator $g$, we have that
$\textup{FID}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  Note that
$\left|\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}-\|\mu_{g}-\mu_{\textup{r}}\|_{2}\right|\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
We derive
@@
$\leq$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
$\leq$ $(2\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.  By Theorem A.3 in Appendix A.5, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
@@
$\left|\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]\right|$
$=$ $\left|\mathbb{E}[\textup{Tr}[(f(X_{g})-\mu_{g})(f(X_{g})-\mu_{g})^{\top}]]-\frac{1}{n}\sum_{i=1}^{n}\textup{Tr}[(f(x^{i})-\widehat{\mu}_{g}^{n})((f(x^{i})-\widehat{\mu}_{g}^{n}))^{\top}]\right|$
$=$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2}\right|$
$\leq$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|.$ (19)
Next, we bound the two terms in the last inequality. Let $Z_{g}:=f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. For the first term, we have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\|Z_{g}\|_{2}^{2}=(\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Sigma_{g}(\Sigma_{g}^{-\frac{1}{2}}Z_{g})=(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Lambda_{g}(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}),$
where we denote by $\Sigma_{g}:=Q_{g}^{\top}\Lambda_{g}Q_{g}$ the eigendecomposition of $\Sigma_{g}$. Note that the random vector $Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}$ has distribution ${\mathcal{N}}(0,I)$. We have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\sum_{j=1}^{d}\lambda_{j}U_{j}^{2},$
where $\lambda_{j}$’s are the eigenvalues of $\Sigma_{g}$, $U_{i}\sim{\mathcal{N}}(0,1)$ are independent standard normal, and hence $U_{i}^{2}$’s are independent Chi-square random variables. Therefore, we have that
$\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|$
$=$ $\left|\sum_{j=1}^{d}\lambda_{j}\left(\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-\mathbb{E}[U_{j}^{2}]\right)\right|$
$\leq$ $\left(\sum_{j=1}^{d}\lambda_{j}\right)\cdot\max_{j=1,\cdots,d}\left|\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-1\right|$
$\leq$ $\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)},$ (20)
where the last inequality holds with probability at least $1-\frac{\delta}{3}$ by Lemma E.5. In addition, for the second term, we derive that
$\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|$
$=$ $\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})(\|f(x^{i})-\mu_{g}\|_{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right|$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right)$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right).$ (21)
@@
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}},$
@@ A.4 Concentration of Mean Vector and Covariance Matrix
Lemma A.2 (Concentration of mean vector and covariance matrix).
Under the same conditions in Theorem A.1, with probability at least $1-\frac{2\delta}{3}$, we have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}},\text{ where }{\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$ (22)
@@
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2},$ (23)
where $\sigma_{g}^{2}:=\max_{i\in[d]}\Sigma_{g}[i,i]$ is the largest variance for the elements in $f(X_{g})$, and $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
@@
1. Concentration of mean vector.  Note that $\Sigma_{g}^{\frac{1}{2}}Z$ where $Z\sim{\mathcal{N}}(0,I)$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution, which is the same as $f(X_{g})-\mu_{g}$. We have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}(\Sigma_{g}^{\frac{1}{2}}Z_{i})\right\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}Z_{i}\right\|_{\Sigma_{g}},$
where $\|y\|_{\Sigma_{g}}:=\sqrt{\langle y,\Sigma_{g}y\rangle}$ is the matrix norm. Let $\widehat{Z}:=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ denote the mean vector. We further obtain that
$\left\|\widehat{Z}\right\|_{\Sigma_{g}}^{2}=\textup{Tr}\left((\widehat{Z})^{\top}\Sigma_{g}\widehat{Z}\right)=\textup{Tr}\left(\Sigma_{g}\widehat{Z}(\widehat{Z})^{\top}\right)=\textup{Tr}(\Sigma_{g}D)+\textup{Tr}(\Sigma_{g}M),$ (24)
where $D:=\textup{diag}(\widehat{Z}(\widehat{Z})^{\top})$ and $M=\widehat{Z}(\widehat{Z})^{\top}-D$. For the first term, note that each diagonal element has the same distribution $\frac{1}{n^{2}}(\sum_{i=1}^{n}z_{i})^{2}$, where $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\frac{\delta}{6}$, we have that
$\textup{Tr}(\Sigma_{g}D)\leq\|D\|_{2}\textup{Tr}(\Sigma_{g})\leq\textup{Tr}(\Sigma_{g})\left(\frac{1}{n}\sqrt{32\log\left(\frac{12d}{\delta}\right)}\right),$ (25)
where the first inequality holds by that $\Sigma_{g}$ is PSD, and the second inequality holds by Lemma E.10. For the second term,
$\textup{Tr}(\Sigma_{g}M)=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot M_{j,l}$
$=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)$
$\leq$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\cdot\max_{j,l}\left|\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)\right|$
where $z_{j,1},\cdots,z_{j,n},z_{l,1},\cdots,z_{l,n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. By Lemma E.11, with probability at least $1-\frac{\delta}{6}$, it holds that
$\textup{Tr}(\Sigma_{g}M)\leq\left(\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\right)\frac{1}{n}\sqrt{8\log\left(\frac{24d}{\delta}\right)}$ (26)
Combining Inequalities (25) and (26), we conclude that with probability at least $1-\frac{\delta}{3}$, it holds that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|^{2}_{2}\leq\left(\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|\right)\left(\frac{2}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)$ (27)
@@
$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2},$ (28)
where the last inequality holds by the triangle inequality of the L2 norm of the matrix. Note that the zero-mean random vector $f(X_{g})-\mu_{g}$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution (under the assumption of FID). For the first term, we evoke Lemma E.7. Then, with probability at least $1-\frac{\delta}{3}$, it holds that
$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}.$ (29)
For the second term, we have that
$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2},$

### v3

@@ A.3 Concentration of Empirical FD (4)
For any generator $g$ and $n\in{\mathbb{N}}_{+}$, we have that
$\textup{FD}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  We derive
@@
$=$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
$\leq$ $(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\text{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.  By Lemma 2, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
@@
$\textup{Tr}[\widehat{\Sigma}_{g}^{n}]=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}\right]$ (21)
$=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right]$
$=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}-(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right]$
$=$ $\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}$
Hence, we obtain
$\left|\textup{Tr}[\Sigma_{g}]-\textup{Tr}[\widehat{\Sigma}_{g}^{n}]\right|\leq\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}$ (22)
Note that $f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. By Lemma 4, with probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\frac{1}{n}\sum_{i=1}^{n}\|f(X^{i})-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]\right|\leq$ $\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$
@@
$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|$
$\leq$ $\underbrace{(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
$+\underbrace{\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right)+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}$
$=$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$
$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
@@ A.4 Concentration of Mean Vector and Covariance Matrix
Lemma 1 ($L$2-norm error for mean and covariance matrix).
Under the same conditions in Theorem 4, with probability at least $1-\frac{2\delta}{3}$, we have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right)}=:\Delta_{\mu_{g}}^{n}$ (23)
@@
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{\mu_{g}}^{n})^{2}=:\Delta_{\Sigma_{g}}^{n}$ (24)
for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
@@
1. Concentration of squared $L$2-norm error.  Note that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}=\underbrace{\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{(*): concentration}}+\underbrace{\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{expected L2-norm error}},$
where $\widehat{\mu}_{g}^{n}-\mu_{g}\sim{\mathcal{N}}(0,\frac{\Sigma_{g}}{n})$. By Lemma 4, we have $\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]=\frac{\textup{Tr}[\Sigma_{g}]}{n}$, and with probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]\right|\leq\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right).$
@@
$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}\left((f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}-2(f(x^{i})-\mu_{g})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}+(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right)\right\|_{2}$
$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}.$ (25)
For the first term, note that $f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. Then, by Lemma 6, with probability at least $1-\frac{\delta}{3}$, it holds that
$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}$ (26)
for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ and $\kappa$ is a constant defined therein. For the second term, we have that
$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}\leq(\Delta_{\mu_{g}}^{n})^{2},$